how do u solve the following:

x=-6

y=6+x

y=-x-1/4

Answer by Fermat(136)

You can't!!
Your equations are inconistent.
If x = -6, then 6+x = 0
Therefore y (=6+x) = 0
But if y=0, then y cannot = -x - 1/4, not if x = -6
The equations are inconistent, therefore there is no solution to the equation set.